OpenStudy (anonymous):

You are a manufacturer of a certain type of pencil. You’ve calculated that it costs $6 to make 300 pencils and $15 to make 1000 pencils. if your Assuming this relationship is linear, write an equation that relates cost to the number of pencils you make.

OpenStudy (anonymous):

-1500 + 700y = 9x 1500 + 700y = 9x 15 + 700y = 9x 1500 + 700y = 11x those are my answers...

OpenStudy (anonymous):

@noniebug how??

OpenStudy (anonymous):

those are four different answers on the worksheet and one of those are the answer.. its asking what form would you put it in.. you solved it.

ganeshie8 (ganeshie8):

-1500 + 700y = 9x